a) Draw the budget constraint for a person with income of $1,000 if the price of Pepsi is $5 and the price of pizza is $10. b) Choose any combination of pizzas and from the table and write a proper budget equation with Pizza denoted as good x and Pepsi denoted as good Y? c) What is the slope of the budget constraint? d) If the person wants to spend equal amount of money on both pizza and Pepsi what combination of quantity of both goods needs to be purchased to satisfy the budget constraint? Write the budget equation in this case? e) What happens to the budget equation and budget set if the income of the person decreases to $500? f) What happens to the budget equation and budget set if the price of pizza decreases to $5 keeping the price of Pepsi and the income same?
Answer #1

a.

Budget is given as:

5 Pepsi + 10 Pizza = 1000

b.

Note: Reference to table missing.

Budget:

10X + 5Y = 1000

c.

d.

To spend equal amount of money on both goods, we have:

10X = 5Y

Using this in budget, we get:

10X + 10X = 1000

X = 50

Y = 100

e.

New budget(BLUE line):

10X + 5Y = 500

Budget line shifts left downwards.

f.

New budget(BLUE line):

5X + 5Y = 1000

Budget pivots and becomes flatter.

If you are satisfied with the answer, please provide a positive rating. Feel free to comment in case of queries.

Have a nice day ahead!